ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 19A22, 20C99, 20B05 Yoshida, Tomoyuki, Hokkaido Univ., Japan Crossed Burnside rings of finite groups Let G be a finite group. A crossed G-set X is a finite G-set equipped with a G-weight function |x|. The Grothendieck ring XΩ(G) of finite crossed Gsets is called the crossed Burnside ring. Then we have an embeding into the crossed ghost ring, an idempotent formula, an action on Mackey funtors.